whats the width of a rectangle that has a perimeter of 36 metres and a length that is three times the width
Answer by macston(5194)

.
W=width; L=length=3W; P=perimeter=36 meters
P=2(L+W)Substitute for L.
36m=2(3W+W) Divide each side by 2.
18m=4W Divide each side by 4
4.5m=W ANSWER: The width is 4.5 meters.
.
CHECK:
.
L=3W=3(4.5m)=13.5m The length is 13.5 meters
P=2(L+W)
36m=2(13.5m+4.5m)
36m=2(18m)
36m=36m